OpenStudy (anonymous):

Jesse's condo has a length 10 feet longer than the width. If his condo is 1,064 square feet, what are the dimensions? A. 28 feet by 38 feet B. 19 feet by 56 feet C. 14 feet by 76 feet D. 32.6 feet by 32.6 feet

OpenStudy (anonymous):

@TechnoSoul @patrickthestar @agent0smith @xxferrocixx @dmytro1996 @emok9893 @MikaylaSkyy @KingGeorge @YINGYANGMU @Elsa213 @RosieF @timo86m @triciaal

OpenStudy (anonymous):

which option has 2 number that have a difference of 10 and multiply to 1064?

OpenStudy (anonymous):

*numbers

OpenStudy (anonymous):

or rather, which has an option where the 2 numbers just have a difference of 10?

OpenStudy (triciaal):

assume condo will be shape of rectangle length l and width w given length 10 feet longer than the width means l = w + 10 area = l*w = 1064

OpenStudy (anonymous):

Okay, so the first thing you need to do is set up the equations the word problem gives you So, you have L = W + 10 -> Because the Length is 10 feet longer the Width and then the equation for area, which is L * W = Area

OpenStudy (triciaal):

@karendiaz123 i will let @MikaylaSkyy continue

OpenStudy (anonymous):

With that, you substitute the equation for L into the equation for Area. So, Area = (W+10) * W And then plug in the area that you know 1064 = (W+10) * W and Solve for W

OpenStudy (anonymous):

Do you know how to do that?